OpenStudy (anonymous):

Trigonometry: How would you transform f(x) = 3 sin(4x - π) + 4 into a cosine function in the form f(x) = a cos(bx - c) + d?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Hero @dan815 @paki

OpenStudy (anonymous):

From my understanding, I'd just have to change sine to cosine and add -pi/2

OpenStudy (anonymous):

So would it just be \[f(x)=3\cos(4x- \pi/2)+4\] ???

OpenStudy (anonymous):

@perl

OpenStudy (michele_laino):

hint: \[\begin{gathered} \sin \left( {4x - \pi } \right) = \sin \left( {4x} \right)\cos \pi - \cos \left( {4x} \right)\sin \pi = \hfill \\ = - \sin \left( {4x} \right) \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

i don't get it

OpenStudy (anonymous):

That doesn't put it into the form f(x) = a cos(bx-c)+d

OpenStudy (michele_laino):

that is the first step

OpenStudy (anonymous):

oh okay thats a trig identity right?

OpenStudy (michele_laino):

yes! It is the subtraction formula for sin function

OpenStudy (anonymous):

sin(α - β) = sinα cos β - cos α sin β ?

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

then what goes after?

OpenStudy (michele_laino):

I'm trying to write the subsequent step, please wait

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

here is the next step: \[\Large \sin \left( {4x} \right) = \cos \left( {4x - \frac{\pi }{2}} \right)\] @genesis98

OpenStudy (michele_laino):

so we can write: \[\Large f\left( x \right) = - 3\cos \left( {4x - \frac{\pi }{2}} \right) + 4\] @genesis98

OpenStudy (anonymous):

oh okay! thank you!

OpenStudy (michele_laino):

thank you!